Determine the resistance and the power absorbed by an electric heater that uses 5.0 A when connected across 110 V.


1
Expert's answer
2021-10-04T04:55:52-0400

Electric heater: 5.0 A

Connected across: 110 V

𝑅𝑅𝑒𝑒𝑒𝑒 = 𝑉𝑉 𝐼𝐼 = 100.0 35 = 20.0 7 Ξ© 1

𝑅𝑅𝑒𝑒𝑒𝑒 = 1 4.0 + 24.0 + 𝑅𝑅 96.0 + 28𝑅𝑅 = 24.0 + 7𝑅𝑅 + 24.0 + 𝑅𝑅 4.0(24.0 + 7𝑅𝑅) = 12.0 + 2𝑅𝑅 24.0 + 7𝑅𝑅 24.0 + 7𝑅𝑅 12.0 + 2𝑅𝑅 = 20.0 7 β†’ 168 + 49𝑅𝑅 = 240 + 40𝑅𝑅 β†’ 9𝑅𝑅 = 72 𝑅𝑅 = 8.0 Ξ©